Net Income Planning Superior Corporation sells a single product for $60 per unit, of which $36 is contribution margin. Fixed costs total $108,000 and net income before income tax is $28,800 Determine the following a. The present sales volume in dollars. $ 0 b. The break-even point in units. 0 units c. The sales volume in units necessary to attain a net income before income tax of $39,600 0 units d. The sales volume in units necessary to attain a net income before income tax equal to 20% of sales revenue. 0 e. The sales volume in units necessary to attain an after-tax net income of $43,200 if the tax rate is 40% units units Check
Answer #1

Answers

A= The present sales volume in $ = $228,000

A

Fixed cost ,

$ 1,08,000.00

B

Operating income

$      28,800.00

C=A+B

Contribution margin'

$ 1,36,800.00

D

Contribution margin per unit

$              36.00

E=C/D

Units sold

                 3,800

F

Sales price per unit

$              60.00

G=E x F

Total sales

$ 2,28,000.00

B Breakeven units = 3000

A

Fixed cost

$ 1,08,000.00

B

Contribution margin per unit

$              36.00

C=A/B

Breakeven units

                 3,000

C Sales to attain netr income of $ 39600 = 4100 units

A

Fixed cost

$ 1,08,000.00

B

Profit required

$      39,600.00

C=A+B

Total Contribution required

$ 1,47,600.00

D

Contribution margin per unit

$              36.00

E=C/D

Breakeven units

                 4,100

D Units to attain 20% profit on sales = 4500 Units

A

Fixed cost

$ 1,08,000.00

B

Profit required

$      54,000.00

C=A+B

Total Contribution required

$ 1,62,000.00

D

Contribution margin per unit

$              36.00

E=C/D

Breakeven units

                 4,500

Current sales

%

Sales revenue

$ 2,70,000.00

100%

Variable cost (60-36)

$ 1,08,000.00

40%

Contribution margin'

$ 1,62,000.00

60%

Fixed cost

$ 1,08,000.00

40%

Operating income before taxes

$      54,000.00

20%

20% profit means the fixed cost must be equal to 40% of sales and then we can get a profit of 20% on sales.

So Sales must be 108000/40%